15. In a multiple dilution series, 10 mcL of serum and 40 mcL of diluent are added and mixed in tube 1. From tube 1, 5 mcL are taken and placed into 20 mcL of diluent in tube 2. What is the final dilution of tube 2? (2 points)
16. Using the same dilution from the previous question, if the original concentration of the sample had been 50 mg/dL, what is the concentration of the sample in tube 1?
1
Expert's answer
2020-04-23T13:10:45-0400

Using dilution formula we can determine the concentration change on each step:

C1V1 = C2V2

tube 1. If the initial concentration was 50 mg/dl, which is 50*10-5 mg/mcl, then at the first step;

50*10-5 mg/mcl * 10 mcl = C2 * (10+40) mcl

C2 = 10*10-5 mg/mcl is the diluted concentration.


In the second tube: 10*10-5 mg/mcl * 1.5 mcl = C2 * 21.5 mcl

C2 = 6.98*10-4 mg/ mcl